40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 8 16:27:13 2009, Jenne, Update, COC, ITMX wiped 
    Reply  Wed Dec 9 18:01:06 2009, Koji, Update, COC, Wiping finished 
Message ID: 2373     Entry time: Wed Dec 9 18:01:06 2009     In reply to: 2367
Author: Koji 
Type: Update 
Category: COC 
Subject: Wiping finished 

[Kiwamu, Jenne, Alberto, Steve, Bob, Koji]

We finished wiping of four test masses without any trouble. ITMY looked little bit dusty, but not as much as ITMX did.
We confirmed the surface of the ITMX again as we worked at vertex a lot today. It still looked clean.

We closed the light doors. The suspensions are left free tonight in order to check their behavior.
Tomorrow morning from 9AM, we will replace the door to the heavy ones.

ELOG V3.1.3-